What are some examples of perpendicular lines and parallel lines of the linear equation: 2x-5y=20 and how do you find/get that answer? Please explain thoroughly and use step by step. Thank you.
Answer by Alan3354(69443)

What are some examples of perpendicular lines and parallel lines of the linear equation: 2x-5y=20 and how do you find/get that answer?
-------------
Find the slope of 2x-5y=20
To find the slope, m, put the equation in slope-intercept form. That means solve for y.
2x-5y=20
-5y = -2x + 20
y = (2/5)x - 4
That's the slope-intercept form, y = mx + b. m is the slope, 2/5
--------------
All lines parallel (all infinity of them) have the same slope.
eg, y = (2/5)x + 100
y = (2/5)x + 1
y = (2/5)x
y = (2/5)x -9/37, etc
y = (2/5)x + b where b is any number makes it parallel.
---------------
Lines perpendicular have a slope that's the negative inverse, = -5/2
y = (-5/2)x + 1
y = (-5/2)x - 1000
y = (-5/2)x + b
All are perpendicular to 2x - 5y = 20